RHIC pp Set-Up: AGS J-10 Dump Bump Power Supply had over temperature alarm, requiring flushing of the cooling lines.
Bad settings from the collimators were discovered (leftover from d-Au running) due to high background rates at both PHENIX and STAR. A new set of collimator settings was loaded giving good background rates in both Blue and Yellow for both PHENIX and STAR.
Much of yesterday was spent working on the rotator ramp. In addition the number of bunches was increased from 56 to 109. On the first attempt at 109x109 with 1.1e11 per bunch, snake losses early in the ramp pulled the permit. A second store with 0.65e11 was lost at store. So a decision was made to run 56x56 at 0.6e11 overnight. Polarimeters measured more 68/58% polarization in the Blue/Yellow. Many problems having to do with hanging the Polarimeter FEC, and operator inexperience made for rough going.
PHENIX wants to take a store to measure the polarization state now that we have gone to radial polarization at 8 oclock. During this store, PHENIX will commission their asymmetry detectors. Comparing to CNI detectors requires a polarization measurement every hour.
Scheduling Physicist: Mike Sivertz | Last updated: |
Email: sivertz@bnl.gov | |
Office Phone: 631-344-6102 | |
Digital Pager: 7528 or 4300 | |
Home Phone: 631-878-9020 |